the sail on the boat is in the shape of a triangle. if the height is 5 times the length of base and the area is 90 square feet, find the dimensions of the sail.
Answer by Cromlix(4381)

Hi there,
Area of triangle = 1/2(base x height)
Make base = x
Height = 5x
Area = 1/2(base x height)
90 = 1/2(x * 5x) (* times)
90 = 1/2(5x^2)
90 = 5x^2/2
Multiply both sides by 2
5x^2 = 180
x^2 = 180/5
x^2 = 36
x = 6
Base = 6m
Height = 30m
Hope this helps :-)